# Insights for Early Dark Energy with Big Bang Nucleosynthesis

### v1

@@ V Discussion and Conclusion
Our results indicate that, when constrained only by helium-4 and deuterium abundances, EDE is generally required to remain subdominant throughout most of the BBN epoch. A narrow window of allowed dominance exists around $3\times 10^{8}$ K to $5\times 10^{8}$ K, and modest spikes are permitted near or just after the deuterium bottleneck. The inclusion of a forecasted 1% helium-3 prior produces minimal change in these constraints, implying that at this level of precision, observations of the primordial helium-3 abundance would not provide much new insight into the early expansion history. When the prior on the baryon-to-photon ratio $\eta_{10}$ is relaxed, EDE is allowed to dominate over a wider range of temperatures; however, the resulting $\eta_{10}$ values are in extreme tension with CMB measurements, emphasizing the importance of complementary cosmological data in constraining EDE.

### v2

@@ V Discussion and Conclusion
Our results indicate that, when constrained only by helium-4 and deuterium abundances, EDE is generally required to remain subdominant throughout most of the BBN epoch. We find a more weakly constrained region around $3\times 10^{8}$ K to $5\times 10^{8}$ K, and modest EDE contributions are permitted near or just after the deuterium bottleneck. However, we note that physically consistent models of EDE are unlikely to allow for spikes of energy density that would saturate the upper bounds in these narrow temperature ranges.
The inclusion of a forecasted 1% helium-3 prior produces minimal change in these constraints, implying that at this level of precision, observations of the primordial helium-3 abundance would not provide much new insight into the early expansion history. When the prior on the baryon-to-photon ratio $\eta_{10}$ is relaxed, significantly larger EDE contributions are allowed throughout BBN; however, the resulting best-fit $\eta_{10}$ values in this scenario are in extreme tension with CMB measurements, emphasizing the importance of complementary cosmological data in constraining EDE.